OpenStudy (anonymous):

2sin(3pie/4) = 1 solve for 0 is less than or equal to x less than 2pie

OpenStudy (anonymous):

\[\sin (3\pi/4)\]=1/2,\[0\le x \le 2 \pi\] where is x in question...?

OpenStudy (anonymous):

what to solve...?

OpenStudy (anonymous):

oh sorry (3pie/4 (x))

OpenStudy (anonymous):

ok...x is in multiplication or division..?

OpenStudy (anonymous):

multiplication

OpenStudy (anonymous):

Oh i solved it!

OpenStudy (aravindg):

wow superb!

hartnn (hartnn):

lol., good:) how many values u got ?

OpenStudy (anonymous):

5

hartnn (hartnn):

:O how ?

OpenStudy (anonymous):

well it says to find within 0- 2pie

OpenStudy (anonymous):

so you have to add the period to the reference angle which is x to solve for the other solutions

OpenStudy (anonymous):

\[\sin ((3 \pi/4 )x)=1/2\] \[(3\pi/4)x=\sin^{-1} (1/2)\] \[(3\pi/4)x=\pi/6,5\pi/6\] \[x=\pi/6*(4/(3\pi)),(5\pi/6)*(4/(3\pi))\]

OpenStudy (anonymous):

Thanks! :)